Given a term in a arithmetic sequence and the common difference find the explicit formula
a12=30 d=6

Answer by jim_thompson5910(35256)

a[n] = a[1] + (n-1)d

a[12] = a[1] + (12-1)*6

30 = a[1] + (12-1)*6

30 = a[1] + 66

30-66 = a[1]

-36 = a[1]

a[1] = -36
-------------------------------------------------------

a[n] = a[1] + (n-1)d

a[n] = -36 + (n-1)6

a[n] = -36 + 6n-6

a[n] = 6n-42


So the explicit formula for the sequence is